a) 0.190g of a gas occupies 250.0ml at STP. what is its molecular height. b) what gas is it?
Answer #1

Ideal gas Lau Pv- nRT Platm, R6-0824 Lat K-m T273K v O.2S0L n= Pv/RT - 1atmo 250L 00821 Latm /kmd 23 0.011/5 Ms mdos= MTdecue